What is volume of a sphere with diameter 18 m or meters. Give your answer in terms of pie please i would really appreciate that
Answer by mananth(12270)

volume of sphere = 4/3 *pi*r^3
=4/3 *pi*9^3
=4/3 *pi*729
=4*243*pi
=972pi